I need the solution set for the following equation; (3t/(t+3)) + 4 = (2/(t+3))
Answer by checkley75(3666)

(3t)/(t+3)+4=2/(t+3) add the two fractions.
(3t-2)/(t+3)=-4
3t-2=-4(t+3)
3t-2=-4t-12
3t+4t=-12+2
7t=-10
t=-10/7 answer.